# Deep Incentive Design with Differentiable Equilibrium Blocks

### 08 Mar 2026 v1

@@ A.2.2 Training pipeline
The target joints are sampled uniformly from the simplex and the noise from a standard normal distribution. The batch size is 64. The network is set up to expect an input shape of $[B,2,16,16,2]$. However by sampling masks uniformly at random, we train the generator to solve the task for games with sizes from $2{\times}2$ up to $16{\times}16$. We run the training with three seeds for $10^{6}$ steps using an exponential optimizer schedule with an ADAM optimizer [Kingma and Ba, 2015] and report the best run. The optimization parameters are $b_{1}=0.9,b_{2}=0.999$ with learning rate $0.01$. We first increase learning rate to $0.01$ over 50000 steps and then start ramping it down over 900000 steps. For each seed training takes about six hours on an H100.
@@ A.3.2 Training Pipeline
The training pipeline is illustrated in Figure˜6. The network has three PO2PO layers with $C_{l}=32$, one PO2O layer and two O2O layers with $C_{l}=32$ nodes each. The batch size $B$ is 64 and we run the training for $2*10^{6}$ steps. We run it with three seeds each for both the $0.01$-ME-CE and $0.01$-ME-CCE and take the best run. The optimization is done with an exponential optimization schedule based on an ADAM optimizer [Kingma and Ba, 2015]. The optimization parameters are $b_{1}=0.9,b_{2}=0.999$ with learning rate $0.01$. We first increase learning rate to $0.01$ over 50000 steps and then start ramping it down over 900000 steps. Moreover, over 500000 steps we increase the penalty weight of the paid contracts (i.e., $\sum_{p}v_{p}(o;\theta,\omega)$) from 0 to $1$ to avoid the network learning zero contracts early on. Each run takes about 15 hours on a NVIDIA A100 GPU. The batched input is set up to be of shape $[B,2,16,16,10,3]$, so the network can handle up to 10 outcomes and $16{\times}16$ action space. However, we sample masks uniformly at random, which are also batched and passed to the network, so that within any batch we train the network for a broad distribution of outcomes between 2 and 10 and action spaces between 2 and 16 actions. The transitions are sampled from a Dirichlet distribution with $\alpha=0.1$. The costs of the agents are sampled from [4,0][-4,0] and the principal payoff is uniformly sampled from [5,5][-5,5].
@@ A.4.2 Training Pipeline
We sample the job time distribution $t_{p,j}\sim{\displaystyle\operatorname{Lognormal}\left(0,0.5\right)}$ . From these we calculate the expected payoffs of the players using the formulas provided in Section˜5.5. The batch size is 64 and we run three seeds and take the best run. The weights are optimized with an exponential optimization schedule based using ADAM [Kingma and Ba, 2015]. The optimization parameters are $b_{1}=0.9,b_{2}=0.999$ with learning rate $0.01$. We first increase learning rate to $0.01$ over 50000 steps and then start ramping it down over 900000 steps. We also ramp up the penalty parameter $\lambda$ from 0 to $0.1$ over 50000 steps. Each complete training run consists of $10^{6}$ training steps and takes roughly 10 hours on a NVIDIA A100 GPU.
@@ A.5 Differentiable Equilibrium Blocks
The games are sampled from the equilibrium invariant embedding [Marris et al., 2023] with their $L_{2}$ norm set to $\sqrt{\prod_{p=1}^{N}|\mathcal{A}_{p}|}$, where $|\mathcal{A}_{p}|$ is the effective action size after masking, to ensure unit variance of all inputs.
@@
  • $A_{p}^{CE}(a_{p}^{\prime},a_{p}^{\prime\prime},a)$ and $A_{p}^{CCE}(a_{p}^{\prime},a)$: The deviation gains, i.e.,
@@ B.1 Discussion
While gradient-based approaches have been tried before in the setting of (ID), the existing works have two important restrictions. First, instead of using $\epsilon$-ME-Eql as solution concept, they generally use the Nash equilibrium and have to make restrictive assumptions on its uniqueness and the monotonicity of the game. Second, the gradient computation is very expensive. Some approaches build on implicit differentiation, which requires inverting a matrix that is quadratic in the game size [Li et al., 2023]. The remaining ones use automatic differentiation which requires running many steps of an iterative solver to converge to an approximate equilibrium and then backpropagating through the whole computational graph. In contrast, deep incentive design computes approximate gradients in time linear in $\mathcal{O}(|\mathcal{A}|)$ by backpropagating through the learned DEB. Moreover, our approach naturally learns to generalize between different contexts, whereas the classical gradient-based MPEC solvers need to be (re)run for a specific context.
@@
is similar in spirit to our approach. It consists of running gradient descent on neural networks to design a game. However, the existing literature has focused on Bayesian games—the natural setting for mechanism design problems. In those games, players are reporting their private types and the famous revelation principle states that any mechanism implementing an outcome in dominant strategies has an outcome-equivalent counterpart where it is a dominant strategy equilibrium for all players to report their types truthfully. Therefore, it is natural to restrict to searching within the space of truthful mechanisms. As the equilibrium is already known for these mechanisms, this circumvents the problem of differentiating through the equilibrium of the resulting game. Our focus is on non-Bayesian normal-form games with discrete action spaces, where truthfulness is not a salient concept as the players do not have private information. Therefore the problem of differentiating through the equilibrium persists and is tackled using our DEBs. We thus consider the existing works in differentiable economics as complementary to deep incentive design. Moreover, in the differentiable economics literature the neural networks have generally been trained as the mechanism themselves, taking as input a distribution of bids and outputting the prices for each problem instance. Each task thus requires a different network and training run. In contrast, our networks are not mechanisms but “mechanism generators”, taking as input the distribution of contexts $\omega\sim\Omega$ and outputting the induced mechanism $G(\theta;\omega)$. These generators are trained once upfront and can be evaluated quickly for a broad class of problems, instead of having to be retrained when facing a different problem instance.

### 12 Mar 2026 v2

@@ A.2.2 Training pipeline
The target joints are sampled uniformly from the simplex and the noise from a standard normal distribution. The batch size is 64. The network is set up to expect an input shape of $[B,2,16,16,2]$. However by sampling masks uniformly at random, we train the generator to solve the task for games with sizes from $2{\times}2$ up to $16{\times}16$. We run the training with three seeds for $10^{6}$ steps using an exponential optimizer schedule with an ADAM optimizer Kingma and Ba [2015] and report the best run. The optimization parameters are $b_{1}=0.9,b_{2}=0.999$ with learning rate $0.01$. We first increase learning rate to $0.01$ over 50000 steps and then start ramping it down over 900000 steps. For each seed training takes about six hours on an H100.
@@ A.3.2 Training Pipeline
The training pipeline is illustrated in Figure˜6. The network has three PO2PO layers with $C_{l}=32$, one PO2O layer and two O2O layers with $C_{l}=32$ nodes each. The batch size $B$ is 64 and we run the training for $2*10^{6}$ steps. We run it with three seeds each for both the $0.01$-ME-CE and $0.01$-ME-CCE and take the best run. The optimization is done with an exponential optimization schedule based on an ADAM optimizer Kingma and Ba [2015]. The optimization parameters are $b_{1}=0.9,b_{2}=0.999$ with learning rate $0.01$. We first increase learning rate to $0.01$ over 50000 steps and then start ramping it down over 900000 steps. Moreover, over 500000 steps we increase the penalty weight of the paid contracts (i.e. $\sum_{p}v_{p}(o;\theta,\omega)$) from 0 to $1$ to avoid the network learning zero contracts early on. Each run takes about 15 hours on a NVIDIA A100 GPU. The batched input is set up to be of shape $[B,2,16,16,10,3]$, so the network can handle up to 10 outcomes and $16{\times}16$ action space. However, we sample masks uniformly at random, which are also batched and passed to the network, so that within any batch we train the network for a broad distribution of outcomes between 2 and 10 and action spaces between 2 and 16 actions. The transitions are sampled from a Dirichlet distribution with $\alpha=0.1$. The costs of the agents are sampled from [4,0][-4,0] and the principal payoff is uniformly sampled from [5,5][-5,5].
@@ A.4.2 Training Pipeline
We sample the job time distribution $t_{p,j}\sim{\displaystyle\operatorname{Lognormal}\left(0,0.5\right)}$ . From these we calculate the expected payoffs of the players using the formulas provided in Section˜5.5. The batch size is 64 and we run three seeds and take the best run. The weights are optimized with an exponential optimization schedule based using ADAM Kingma and Ba [2015]. The optimization parameters are $b_{1}=0.9,b_{2}=0.999$ with learning rate $0.01$. We first increase learning rate to $0.01$ over 50000 steps and then start ramping it down over 900000 steps. We also ramp up the penalty parameter $\lambda$ from 0 to $0.1$ over 50000 steps. Each complete training run consists of $10^{6}$ training steps and takes roughly 10 hours on a NVIDIA A100 GPU.
@@ A.5 Differentiable Equilibrium Blocks
The games are sampled from the equilibrium invariant embedding Marris et al. [2023] with their $L_{2}$ norm set to $\sqrt{\prod_{p=1}^{N}|\mathcal{A}_{p}|}$, where $|\mathcal{A}_{p}|$ is the effective action size after masking, to ensure unit variance of all inputs.
@@
  • $A_{p}^{CE}(a_{p}^{\prime},a_{p}^{\prime\prime},a)$ and $A_{p}^{CCE}(a_{p}^{\prime},a)$: The deviation gains, i.e.
@@ B.1 Discussion
While gradient-based approaches have been tried before in the setting of (ID), the existing works have two important restrictions. First, instead of using $\epsilon$-ME-Eql as solution concept, they generally use the Nash equilibrium and have to make restrictive assumptions on its uniqueness and the monotonicity of the game. Second, the gradient computation is very expensive. Some approaches build on implicit differentiation, which requires inverting a matrix that is quadratic in the game size Li et al. [2023]. The remaining ones use automatic differentiation which requires running many steps of an iterative solver to converge to an approximate equilibrium and then backpropagating through the whole computational graph. In contrast, deep incentive design computes approximate gradients in time linear in $\mathcal{O}(|\mathcal{A}|)$ by backpropagating through the learned DEB. Moreover, our approach naturally learns to generalize between different contexts, whereas the classical gradient-based MPEC solvers need to be (re)run for a specific context.
@@
is similar in spirit to our approach. It consists of running gradient descent on neural networks to design a game. However, the existing literature has focused on Bayesian games—the natural setting for mechanism design problems. In those games, players are reporting their private types and the famous revelation principle states that any mechanism implementing an outcome in dominant strategies has an outcome-equivalent counterpart where it is a dominant strategy equilibrium for all players to report their types truthfully. Therefore, it is natural to restrict to searching within the space of truthful mechanisms. As the equilibrium is already known for these mechanisms, this circumvents the problem of differentiating through the equilibrium of the resulting game. Our focus is on non-Bayesian normal-form games with discrete action spaces, where truthfulness is not a salient concept as the players do not have private information. Therefore the problem of differentiating through the equilibrium persists and is tackled using our DEBs. We thus consider the existing works in differentiable economics as complementary to deep incentive design. Moreover, in the differentiable economics literature the neural networks have generally been trained as the mechanism themselves, taking as input a distribution of bids and outputting the prices for each problem instance. Each task thus requires a different network and training run. In contrast our networks are not mechanisms but “mechanism generators”, taking as input the distribution of contexts $\omega\sim\Omega$ and outputting the induced mechanism $G(\theta;\omega)$. These generators are trained once upfront and can be evaluated quickly for a broad class of problems, instead of having to be retrained when facing a different problem instance.